# Beyond LLMs: An Exploration of Small Open-source Language Models in Logging Statement Generation

### v1

@@ Abstract.
Effective software maintenance heavily relies on high-quality logging statements, but manual logging is challenging, error-prone, and insufficiently standardized, often leading to inconsistent log quality. While large language models have shown promise in automatic logging, they introduce concerns regarding privacy, resource intensity, and adaptability to specific enterprise needs. To tackle these limitations, this paper empirically investigates whether Small Open-source Language Models (SOLMs) could become a viable alternative via proper exploitation. Specifically, we conduct a large-scale empirical study on four prominent SOLMs, systematically evaluating the impacts of various interaction strategies, parameter-efficient fine-tuning techniques, model sizes, and model types in automatic logging. Our key findings reveal that Retrieval-Augmented Generation significantly enhances performance, and LoRA is a highly effective PEFT technique. While larger SOLMs tend to perform better, this involves a trade-off with computational resources, and instruct-tuned SOLMs generally surpass their base counterparts. Notably, fine-tuned SOLMs, particularly Qwen2.5-coder-14B, outperformed existing specialized tools and LLM baselines in accurately predicting logging locations and generating high-quality statements, a conclusion supported by traditional evaluation metrics and LLM-as-a-judge evaluations. Furthermore, SOLMs also demonstrated robust generalization across diverse, unseen code repositories.
Logging Statement, Logging Practice, Large Language Model
@@ 1. Introduction
Logs are textual records generated during software execution to capture runtime events, states, and contextual information (Yu et al., 2023). A typical log statement consists of three components: a verbosity level, logging variables, and logging texts (He et al., 2021). In particular, as demonstrated in the example below, the logging level (such as debug) reflects the event’s severity; the logging variables (like terminalState) hold critical run-time data about system states; meanwhile, the logging text (for instance, Stopping the checkpoint services with state) offers a static explanation of the system’s actions. High-quality logs provide actionable insights for developers to diagnose failures, optimize system behavior, and ensure operational reliability. However, the absence or inadequacy of logging statements can severely hinder downstream tasks such as anomaly detection and failure diagnosis, leading to prolonged debugging cycles and increased maintenance costs. Consequently, the strategic placement and content design of log statements directly influence the effectiveness of software maintenance and evolution (Soldani and Brogi, 2022).
@@
Manual logging poses significant challenges for developers, making it a complex and error-prone task. First, the absence of universal logging guidelines results in inconsistent practices across development teams, where decisions about log placement, content, and verbosity heavily rely on individual developers’ expertise and intuition. This lack of standardization often leads to logs that vary widely in quality, granularity, and utility (Chen and Jiang, 2019). Second, striking a balance between log quantity and quality is inherently difficult: excessive logging generates vast amounts of data, increasing storage costs, slowing down system performance, and overwhelming analysis tools with irrelevant information, while insufficient logging risks omitting critical runtime details necessary for diagnosing failures or understanding system behavior (Yao et al., 2018). Moreover, manual logging is time-intensive, requiring developers to anticipate potential failure points and system states during code writing, which is challenging in large, complex, or rapidly evolving software systems. The cognitive burden of determining optimal logging points and crafting meaningful log messages further exacerbates the problem, often leading to logs that are either too vague to be actionable or overly detailed, obscuring key insights (Hassani et al., 2018). Additionally, as software evolves, manually inserted logs may become outdated or misaligned with new system behaviors, requiring continuous maintenance to remain relevant. These challenges highlight the need for automated, systematic approaches to logging (Zhong et al., 2025).
@@
To address these challenges, at the early stage, researchers have explored automated logging techniques by decomposing the problem into sub-tasks, such as where-to-log (identifying code locations for logging) (Zhao et al., 2017; Li et al., 2020), what-to-log (generating static content) (Ding et al., 2022; Liu et al., 2019), and log-level suggestion (Heng et al., 2024; Li et al., 2021a; Liu et al., 2022). However, these fragmented approaches lack integration into a unified, end-to-end framework for generating complete logging statements. Recent advances in pre-trained language models (LMs) have opened new avenues for automated logging. LANCE (Mastropaolo et al., 2022) and LEONID (Mastropaolo et al., 2024) pioneered the use of sequence-to-sequence models (e.g., T5 (Raffel et al., 2020)) to generate logging statements directly from code contexts. Subsequent tools, such as, Fastlog (Xie et al., 2024), Unilog (Xu et al., 2024), and SCLogger (Li et al., 2024b), further leveraged large language model (LLM) to improve logging quality. Notably, Unilog and SCLogger adopted prompt-based methods with LLMs like GPT-3.5 and Codex, achieving state-of-the-art performance by harnessing LLMs’ code comprehension and natural language generation capabilities.
Despite their effectiveness, LLM-based logging tools face the following three limitations. First, privacy concerns arise when proprietary code is processed via commercial LLM APIs (e.g., OpenAI), risking exposure of sensitive intellectual property (Yao et al., 2024). For instance, Samsung banned employee use of ChatGPT and other generative AI tools after an engineer accidentally leaked sensitive source code to ChatGPT (Kharpal, 2023). Second, the resource intensity of LLMs—requiring thousands of GPU hours for training and generating substantial carbon emissions—contradicts sustainable computing goals (Imani et al., 2024). Researchers have shown that making requests to LLMs can lead to annual carbon emissions greater than the emission during training such LLMs (Chien et al., 2023). Additionally, LLMs often often struggle to adapt to enterprise-specific logging styles due to varying organizational needs (He et al., 2022; Gu et al., 2022; Rong et al., 2020). Enterprises may require unique verbosity levels, from concise logs for performance to detailed ones for debugging, which LLMs may not consistently deliver. Error prioritization also differs, with some organizations focusing on critical alerts and others on detailed behavioral logs, challenging LLMs’ generic outputs. Compliance-driven requirements, such as general data protection regulation (European Parliament and Council of the European Union, 2016), demand specific log formats or data anonymization that LLMs may not inherently support. This misalignment often necessitates manual refinement of LLM-generated logs to meet enterprise standards.
Recently, Small open-source language model (SOLM) has demonstrated competitive performance in many software engineering tasks (Lu et al., 2023; Imani et al., 2024), such as program repair (Silva et al., 2023) and comment rectification (Rong et al., 2024). Thus, SOLMs, which defined as open-source models smaller than 14B parameters, offer a promising alternative to LLMs because it can theoretically address the above three limitations in logging automation. By enabling local deployment on a single consumer-grade graphics card (e.g., A100) (Yang et al., 2024b), SOLMs mitigate the privacy risks associated with processing proprietary code through commercial LLM APIs. Additionally, their reduced computational requirements lower resource costs and carbon emissions, aligning with sustainable computing goals. Furthermore, SOLMs can be fine-tuned on enterprise codebases to adapt to unique logging styles, addressing the challenge of aligning logs with organizational practices. However, despite their theoretical ability, no empirical studies have systematically evaluated SOLMs’ feasibility for this task.
@@
In this paper, we conduct an empirical study on four prominent SOLMs, namely LLaMA, Mistral, CodeLLaMA, and Qwen2.5coder, to explore the potential utility of SOLMs in the automatic generation of logging statements. For a thorough evaluation of the effectiveness of SOLMs, we pose the following four research questions:
@@
RQ1: What are the most effective interaction strategies for using SOLMs in logging generation? Diverse prompting techniques can influence the performance of SOLMs without the need for retraining. Gaining insight into their effects can help improve the generation of logging statements across different contexts.
@@
Result. Retrieval-augmented generation outperforms in-context learning and chain-of-thought, significantly enhancing the performance of logging automation task.
@@
RQ2: What’s the best strategy using SOLMs for automated logging statement generation? Many strategies such as parameter-efficient fine-tuning (PEFT) techniques, model size, and model type may influence the efficacy. Thus, we further investigate the extent of their impact, which may offer insights into the optimal selection of strategies for enhancing SOLM peformance.
@@
Result. LoRA demonstrate the most consistent and superior results when fine-tuning with PEFT techniques. For models with more than 3B parameters, performance in generating logging statement improves with more parameters, but increases computational costs, indicating a performance-resource trade-off. The instruct variant of SOLM model outperforms its base counterpart, benefiting from its instruction-tuned foundation.
@@
RQ3: How effectively do SOLMs compare to existing methods and LLM baselines in automated logging statement generation? After eastablishing the optimal strategies for employing SOLMs, we aims to investigate the performance of SOLMs in automated logging compared to existing method and the prominent LLMs.
@@
Result. Fine-tuned SOLM, particularly Qwen2.5-coder-14B, outperform both existing methods and LLMs across all evaluated metrics, demonstrating superior logging location accuracy and statement quality. The result of LLM judger further supports the high quality of SOLM-generated statements.
@@
Result. SOLMs demonstrate strong generalization capabilities in automated logging statement generation, maintaining high performance across diverse, unseen repositories. Similar logging practices, such as those shared among Apache open-source projects, significantly improve the cross-project generalization ability of SOLMs.
@@
  • We conduct the first large-scale empirical study to investigate the capability of SOLMs for automated logging statement generation, which demonstrating their capability to produce contextually and syntactically correct logging statement comparable to existing methods and LLMs.
  • We explore parameter-efficient fine-tuning techniques that enhance SOLM performance while largely reducing computational cost, thus enabling effective fine-tuning on limited resources.
  • We demonstrate SOLMs’ generalization and practical advantages. SOLMs exhibit strong generalization across diverse, unseen code repositories, while their local deployment capability ensures privacy and aligns with sustainable computing goals.
  • We make the source code, dataset, and results of our study publicly available (Renyi Zhong, Yichen Li, Guangba Yu, Wenwei Gu, Jinxi Kuang, Yintong Huo, Michael R. Lyu, 2025) to benefit both practitioners and researchers in the field of automated logging statement generation.
Paper Orgenizations. Section 2 discusses the background. Section 3 describes the experimental design of our study. Section 4 presents the experimental results. Section 5 introduces the advantages of using SOLMs for automated logging statement generation and potential future work directions. Section 6 discusses threats to validity. Section 7 introduces the related work. Section 8 concludes the paper.
@@ 2.1. Problem Definition
This paper focuses on the automated logging statement generation task (i.e., where-to-log + what-to-log), which to some extent can be viewed as a code editing problem: when presented with lines of code, usually corresponding to a method, the generator’s task is to identify both the precise location for logging, referred to as the logging point, and the complete logging statement (i.e., level, variables, and text). The predicted logging point should match the one that was originally present in the source file before being removed, and the predicted logging statement itself should closely resemble the excised original. Figure 1 provides a visual example of this task, showing how a proficient logging statement generator would intelligently incorporate LOG.info("Generating " + totalRows + " using " + numSplits); at line 4. It is important to highlight that this task is distinctly separate from the comprehensive empirical investigation conducted by Li et al. (Li et al., 2024a), which predominantly examines the question of what-to-log but lacks the consideration of where-to-log.
@@ 2.3. LLM Applications in Software Engineering Task
In certain tasks, approaches that utilize prompts have been shown to yield superior outcomes (Imani et al., 2024). Pan et al., for instance, studied the efficiency of LLMs in the context of code translation (Pan et al., 2024). Within the range of models assessed, which encompassed both SOLMs and GPT-4, the top-performing SOLM, known as StarCoder, managed to reach a successful translation rate of 14.5%, in contrast to the success rate of 47.3% achieved by GPT-4. In contrast, SOLMs have been shown to achieve comparable outcomes in specific domains. For instance, Tian et al. (Tian et al., 2024) observed an F1-score of 86.58% using UniXCoder for the task of detecting equivalent mutants, providing a notable contrast to the performance of GPT-4, which recorded an F1-score of approximately 55.90%. Furthermore, when employing Vicuna 13B in conjunction with the innovative LogPrompt strategy, the performance was found to be on par with that of GPT-4, as reported in (Liu et al., 2024b).
@@ 3. Experimental Design
Figure 2 illustrates the overview of our study design. Initially, based on the AL-Bench dataset (Tan et al., [n. d.]), we construct the fine-tuning, valid and test datasets. Then we investigate SOLMs for automated logging statement generation through four research questions. RQ1 seek to identify the most effective interaction strategies for using SOLMs in this task. RQ2 aims to determine the optimal fine-tuning strategies by PEFT techniques, model sizes, and model types. Following that, RQ3 compares the performance of fine-tuned SOLMs against existing specialized methods and LLM baselines. Finally, RQ4 assesses the ability of SOLMs to generalize their logging statement generation capabilities across diverse and unseen code repositories.
@@ 3.1. Dataset Preparation
To evaluate the performance of automated logging statement generation, we select the most recently proposed AL-Bench (Tan et al., [n. d.]). This dataset, containing 42,224 logging statements from 10 widely used projects, was specifically constructed using stringent criteria (e.g., ¿=10k stars, ¿=1k logging statements, ¿=500 log-related issues per project) to ensure the inclusion of well-maintained repositories where developers prioritize high-quality logging statements. Furthermore, AL-Bench intentionally incorporates projects from diverse domains—such as database management, task scheduling, distributed systems, messaging, and IoT platforms—to ensure comprehensive coverage of various real-world logging requirements and practices, making it a robust benchmark. Measures like using recent project versions and standardized code formatting were also employed during its creation to mitigate potential data contamination risks from pre-training corpora. Table 1 provides further statistical details on these source repositories.
@@
To align with the research objectives of this study and accommodate the requirements of certain baseline models, we performed several adjustments to the original dataset. Firstly, recognizing that some baseline models have an input token limit, we filtered out data instances where the input code exceeded 512 tokens. Secondly, the original construction of AL-Bench could generate multiple data points from a single Java function if it contained multiple logging statements, with each data point representing one specific automated logging statement generation case. To prevent potential data leakage, where highly similar code snippets from the same source file might inadvertently appear across different dataset splits (e.g., fine-tuning and testing), we implemented a file-level splitting strategy. This approach ensures that all data instances originating from the same Java file are strictly allocated to only one of the fine-tuning, validation, or test sets. After applying these pre-processing steps, we obtained a final dataset comprising 33,224 instances. We then partitioned this dataset into fine-tuning, validation, and test sets, targeting an 8:1:1 ratio. Because our file-level splitting strategy required keeping all instances from a single file within the same set, the resulting distribution was approximate. The final fine-tuning set contains 26,713 instances, the validation set contains 3,508 instances, and the test set contains 3,003 instances.

### v2

# Larger Is Not Always Better: Exploring Small Open-source Language Models in Logging Statement Generation

@@ Abstract.
Developers use logging statements to create logs that document system behavior and aid in software maintenance. As such, high-quality logging is essential for effective maintenance; however, manual logging often leads to errors and inconsistency. Recent methods emphasize using large language models (LLMs) for automated logging statement generation, but these present privacy and resource issues, hindering their suitability for enterprise use. This paper presents the first large-scale empirical study evaluating small open-source language models (SOLMs) for automated logging statement generation. We evaluate four prominent SOLMs using various prompt strategies and parameter-efficient fine-tuning techniques, such as Low-Rank Adaptation (LoRA) and Retrieval-Augmented Generation (RAG). Our results show that fine-tuned SOLMs with LoRA and RAG prompts, particularly Qwen2.5-coder-14B, outperform existing tools and LLM baselines in predicting logging locations and generating high-quality statements, with robust generalization across diverse repositories. These findings highlight SOLMs as a privacy-preserving, efficient alternative for automated logging.
Software Logging, Logging Statement, Logging Text, Logging Practice, Large Language Model
@@ 1. Introduction
Logs are textual records generated during software execution to capture runtime events, states, and contextual information (Yu et al., 2023b). A typical log statement consists of three components: a verbosity level, logging variables, and logging texts (He et al., 2021; Jiang et al., 2024b). In particular, as the example shown below, the logging level (e.g., debug) reflects the event’s severity; the logging variables (e.g., terminalState) hold critical run-time data about system states; meanwhile, the logging text (e.g., Stopping the checkpoint services with state) offers a static explanation of the system’s actions.
@@
High-quality logs provide actionable insights for developers to diagnose failures, optimize system behavior, and ensure reliability. However, the absence or inadequacy of logging statements can severely hinder downstream tasks such as anomaly detection (Zhang et al., 2022; Li et al., 2020b; Liu et al., 2023) and failure diagnosis (Yu et al., 2023a; Jiang et al., 2025; Huang et al., 2024b), leading to prolonged debugging cycles and increased maintenance costs. Consequently, the strategic placement and content of log statements directly influence the effectiveness of software maintenance and evolution (Soldani and Brogi, 2022; Huang et al., 2024a).
Despite their critical role in software maintenance, producing high-quality logs manually is far from straightforward for developers. First, the absence of universal logging guidelines leads to inconsistent practices, where log quality, granularity, and utility vary widely based on individual developers’ expertise. This inconsistency complicates log analysis and reduces their diagnostic value (Chen and Jiang, 2019). Second, developers face the difficult task of balancing log quantity and quality: over-logging burdens systems with excessive data, while under-logging risks missing critical information (Yao et al., 2018). Third, the cognitive and time-intensive nature of manual logging further exacerbates these issues, as developers must anticipate complex system behaviors and failure points, often resulting in logs that are either too vague or overly detailed (Hassani et al., 2018). Additionally, maintaining log relevance over time is challenging, as software evolution can render existing logs obsolete (Zhong et al., 2025). These challenges highlight the need for automated logging statement generation (hereafter referred to simply as ‘automated logging’) solutions that can consistently produce high-quality logs.
@@
To address these challenges, at the early stage, researchers have explored automated logging techniques by decomposing the problem into sub-tasks, such as where-to-log (identifying code locations for logging) (Zhao et al., 2017; Li et al., 2020a), what-to-log (generating static content and dynamic variables) (Ding et al., 2022; Liu et al., 2019), and log-level suggestion (Heng et al., 2024; Li et al., 2021a; Liu et al., 2022). However, these fragmented approaches lack integration into a unified, end-to-end framework for generating complete logging statements. Recent advances in pre-trained language models (LMs) have opened new avenues for automated logging. LANCE (Mastropaolo et al., 2022) and LEONID (Mastropaolo et al., 2024) pioneered the use of sequence-to-sequence models (e.g., T5 (Raffel et al., 2020)) to generate logging statements directly from code contexts. Subsequent tools, such as Fastlog (Xie et al., 2024), Unilog (Xu et al., 2024), and SCLogger (Li et al., 2024b), further leveraged the large language model (LLM) to improve logging quality. In particular, Unilog and SCLogger adopted prompt-based methods with LLMs such as GPT-3.5 and Codex, achieving state-of-the-art performance by harnessing the code comprehension and natural language generation capabilities of LLMs.
Despite their effectiveness, LLM-based logging tools (Xie et al., 2024; Xu et al., 2024; Li et al., 2024b) face several limitations in enterprise settings.
  • Privacy Risks. Sending proprietary code to commercial LLM APIs, such as OpenAI’s, risks exposing sensitive intellectual property (Yao et al., 2024). For instance, Samsung banned employee use of ChatGPT and other generative AI tools after an engineer accidentally leaked sensitive source code to ChatGPT (Kharpal, 2023).
  • Style Misalignment. LLMs, trained on generic datasets, struggle to generate logs that align with enterprise-specific logging styles, such as unique verbosity levels or error prioritization requirements, limiting their utility for organizational needs (He et al., 2022; Gu et al., 2022).
To address these issues, enterprises often consider fine-tuning and deploying open-source LLMs in private environments. However, this approach demands substantial computational resources, including thousands of GPU hours, which is impractical for many resource-constrained organizations (Chien et al., 2023; Imani et al., 2024). These limitations necessitate more accessible and efficient solutions for automated logging.
Small open-source language models (SOLMs), defined as open-source models with fewer than 14 billion parameters, have gained traction in software engineering tasks, such as program repair (Silva et al., 2023) and comment rectification (Rong et al., 2024), making them a promising solution for automated logging. By deploying SOLMs locally on consumer-grade hardware, such as an A100 GPU, enterprises can safeguard proprietary code, eliminating privacy risks associated with commercial APIs (Yang et al., 2024b). Moreover, SOLMs require significantly fewer computational resources, reducing costs and aligning with sustainable computing goals (Imani et al., 2024). Additionally, SOLMs can be fine-tuned on enterprise-specific codebases to produce logs that meet unique organizational requirements, such as specific formats or compliance standards (Lu et al., 2023; Rong et al., 2024). For resource-constrained enterprises, SOLMs offer a practical balance of privacy, efficiency, and adaptability, making their application to automated logging highly attractive. However, to the best of our knowledge, no studies have systematically explored the effectiveness of SOLMs in automated logging.
@@
To fill this significant gap, in this paper, we conduct an empirical study on four prominent SOLMs, namely LLaMA (Grattafiori et al., 2024), Mistral (Jiang et al., 2023), CodeLLaMA (Roziere et al., 2023), and Qwen2.5coder (Hui et al., 2024), to explore the potential utility of SOLMs in automatic generation of logging statements. We pose four research questions to comprehensively assess the potential of SOLMs.
@@
RQ1: What are the most effective prompting strategies for using SOLMs in logging generation? Different prompting techniques (e.g., in-context learning (ICL) (Brown et al., 2020), chain-of-thought (COT) (Wei et al., 2022), retrieval-augmented generation (RAG) (Lewis et al., 2020)) can influence the performance of SOLMs without the need for retraining. Gaining insight into their effects can help improve the generation of logging statements across different contexts.
@@
Result. RAG outperforms ICL and COT, significantly enhancing the performance of logging automation task.
@@
RQ2: What is the best tuning strategy using SOLMs for automated logging? Many strategies such as parameter-efficient fine-tuning (PEFT) techniques (Houlsby et al., 2019; Hu et al., 2022), model size, and model type may influence the efficacy. Thus, we further investigate the extent of their impact, which may offer insights into the optimal selection of strategies for enhancing SOLM performance.
@@
Result. LoRA (Hu et al., 2022) demonstrates the most consistent and superior results when fine-tuning with PEFT techniques. For models with more than 3B parameters, performance in generating logging statements improves with more parameters, but the increased computational costs indicate a trade-off between performance and resource costs. The instruct variant of the SOLM model outperforms its base counterpart, benefiting from its instruction-tuned foundation.
@@
RQ3: How effectively do SOLMs compare to existing methods and LLM baselines in automated logging? Upon recognizing the optimal strategies for employing SOLMs, we aim to investigate the performance of SOLMs in automated logging compared to existing methods and the prominent LLMs.
@@
Result. Fine-tuned SOLMs, particularly Qwen2.5-coder-14B, outperform both existing methods and LLMs across all evaluated metrics, demonstrating superior logging location accuracy and statement quality. The result of the LLM-based judger further supports the high quality of SOLM-generated statements.
@@
Result. SOLMs demonstrate strong generalization capabilities in automated logging, maintaining high performance across diverse, unseen repositories. Similar logging practices, such as those shared among Apache open-source projects, significantly improve the cross-project generalization ability of SOLMs.
@@
  • We conduct the first large-scale empirical study assessing the effectiveness of SOLMs for automated logging. Our findings demonstrate their capability to produce contextually accurate and syntactically correct logging statements that rival or exceed the performance of existing specialized methods and LLMs.
  • We investigate and identify effective strategies for optimizing SOLM performance in the context of automated logging. This includes demonstrating that RAG significantly enhances performance and that Low-Rank Adaptation stands out as a highly effective PEFT technique.
  • We showcase the practical advantages of fine-tuned SOLMs, including their robust generalization capabilities across diverse and previously unseen code repositories. This research highlights that SOLMs can maintain high performance in varied settings and offer benefits such as local deployment for data privacy and alignment with enterprise-specific logging styles.
  • To facilitate further research and encourage practical adoption in the field of automated logging, we publicly release our source code, datasets, and comprehensive experimental results (Renyi Zhong, Yichen Li, Guangba Yu, Wenwei Gu, Jinxi Kuang, Yintong Huo, Michael R. Lyu, 2025).
Paper Orgenizations. Section 2 discusses the background. Section 3 describes the experimental design of our study. Section 4 presents the experimental results. Section 5 introduces the advantages of using SOLMs for automated logging and potential future work directions. Section 6 discusses threats to validity. Section 7 introduces the related work. Section 8 concludes the paper.
@@ 2.1. Problem Definition
This paper focuses on the automated logging task (i.e., where-to-log + what-to-log), which to some extent can be viewed as a code editing problem: when presented with lines of code, usually corresponding to a method, the generator’s task is to identify both the precise location for logging, referred to as the logging point, and the complete logging statement (i.e., level, variables, and text). The predicted logging point should match the one that was originally present in the source file before being removed, and the predicted logging statement itself should closely resemble the excised original. Figure 1 provides a visual example of this task, showing how a proficient logging statement generator would intelligently incorporate LOG.info("Generating " + totalRows + " using " + numSplits); at line 4. It is important to highlight that this task is distinctly separate from the comprehensive empirical investigation conducted by Li et al. (Li et al., 2024a), which predominantly examines the question of what-to-log but lacks the consideration of where-to-log.
@@ 2.3. LLM Applications in Software Engineering Task
In certain tasks, approaches that utilize prompts have been shown to yield superior outcomes (Imani et al., 2024; Jiang et al., 2024a). Pan et al., for instance, studied the efficiency of LLMs in the context of code translation (Pan et al., 2024). Within the range of models assessed, which encompassed both SOLMs and GPT-4, the top-performing SOLM, known as StarCoder, managed to reach a successful translation rate of 14.5%, in contrast to the success rate of 47.3% achieved by GPT-4. In contrast, SOLMs have been shown to achieve comparable outcomes in specific domains. For instance, Tian et al. (Tian et al., 2024) observed an F1-score of 86.58% using UniXCoder for the task of detecting equivalent mutants, providing a notable contrast to the performance of GPT-4, which recorded an F1-score of approximately 55.90%. Furthermore, when employing Vicuna 13B in conjunction with the innovative LogPrompt strategy, the performance was found to be on par with that of GPT-4, as reported in (Liu et al., 2024b).
@@ 3. Experimental Design
Figure 2 illustrates the overview of our study design. Initially, based on the AL-Bench dataset (Tan et al., [n. d.]), we construct the fine-tuning, valid and test datasets. Then we investigate SOLMs for automated logging through four research questions (RQs). RQ1 investegates the most effective prompt strategies for using SOLMs in this task. RQ2 aims to determine the optimal fine-tuning strategies by PEFT techniques, model sizes, and model types. Following that, RQ3 compares the performance of fine-tuned SOLMs against existing methods and LLM baselines. Finally, RQ4 assesses the ability of SOLMs to generalize their logging statement generation capabilities across diverse unseen code repositories.
@@ 3.1. Dataset Preparation
To evaluate the performance of automated logging, we select the most recently proposed AL-Bench (Tan et al., [n. d.]). This dataset, containing 42,224 logging statements from 10 widely used projects, was specifically constructed using stringent criteria (e.g., ¿=10k stars, ¿=1k logging statements, ¿=500 log-related issues per project) to ensure the inclusion of well-maintained repositories where developers prioritize high-quality logging statements. Furthermore, AL-Bench intentionally incorporates projects from diverse domains, such as database management, task scheduling, distributed systems, messaging, and IoT platforms, to ensure comprehensive coverage of various real-world logging requirements and practices, making it a robust benchmark. Measures like using recent project versions and standardized code formatting were also employed during its creation to mitigate potential data contamination risks from pre-training corpora. Table 1 provides further statistical details on these source repositories.
@@
To align with the research objectives of this study and accommodate the requirements of certain baseline models, we performed several adjustments to the original dataset. Firstly, recognizing that some baseline models have an input token limit, we filtered out data instances where the input code exceeded 512 tokens. Secondly, the original construction of AL-Bench could generate multiple data points from a single Java function if it contained multiple logging statements, with each data point representing one specific automated logging case. To prevent potential data leakage, where highly similar code snippets from the same source file might inadvertently appear across different dataset splits (e.g., fine-tuning and testing), we implemented a file-level splitting strategy. This approach ensures that all data instances originating from the same Java file are strictly allocated to only one of the fine-tuning, validation, or test sets. After applying these pre-processing steps, we obtained a final dataset comprising 33,224 instances. We then partitioned this dataset into fine-tuning, validation, and test sets, targeting an 8:1:1 ratio. Because our file-level splitting strategy required keeping all instances from a single file within the same set, the resulting distribution was approximate. The final fine-tuning set contains 26,713 instances, the validation set contains 3,508 instances, and the test set contains 3,003 instances.